# Mind the Gap: Pitfalls of LLM Alignment with Asian Public Opinion

### 06 Mar 2026 v1

@@ Introduction
Large language models (LLMs) have become essential tools for accessing information and generating content, with platforms such as ChatGPT handling billions of prompts from a global user base (Backlinko 2025; TechCrunch 2025). As of December 2025, ChatGPT was ranked 5th among the world’s most visited websites according to Similarweb (Similarweb 2025). Additionally, on social media platforms such as LinkedIn, recent surveys suggest that over 50% of long-form posts may be written or influenced by generative AI tools (Elad 2025). Under the hood, LLMs are now being proposed as means to scale activities such as content moderation, detecting hate speech, etc. (Kumar et al. 2024). However, this widespread adoption comes with critical challenges. The probabilistic nature of LLMs leads to models preferentially generating viewpoints that are highly represented, and consequently, a biased world-view is derived from its training corpus (Bender et al. 2021a; Seth et al. 2025). Since Internet corpora are heavily skewed toward English, models may disproportionately reflect Western cultural sensibilities (Joshi et al. 2020). This risks marginalizing non-Western perspectives and also may lead to the dissemination of harmful stereotypes, such as linking specific religions to violence (Abid et al. 2021). As these models are increasingly integrated into education, research, and other everyday tasks, their potential to shape public discourse in ways that reinforce existing prejudices becomes a significant concern (Weidinger et al. 2021; Bender et al. 2021b; Santurkar et al. 2023).
@@
While efforts to mitigate these linguistic and cultural biases are ongoing, research on cultural alignment has largely centred on American citizens and has been conducted almost exclusively in English (Santurkar et al. 2023; Durmus et al. 2023). This approach not only overlooks the majority of the world’s population but also ignores the fact that an LLM’s responses can vary significantly depending on the language of the prompt (Kang and Kim 2025). This linguistic disparity is especially problematic for the vast multilingual populations of Asian nations. For instance, while religion’s role has declined in many Western nations (Pew Research Center 2025, 2018; Australian Bureau of Statistics 2022), it remains a central and politically significant aspect of society across much of Asia (The Print 2025; Elvia Muthiariny 2024; Pew Research Center 2023). For billions of multilingual and non-English speaking users, ensuring that LLMs are culturally and linguistically representative is a critical challenge that must be addressed (Santurkar et al. 2023; Green and others 2023).
@@
Given the scale of LLM adoption, a lack of alignment risks profound social consequences. These risks are already evident on social media, where the proliferation of AI-generated content has contributed to polarised discourse and marginalised certain groups, such as the LGBTQ+ community (Kerwin 2024; Bakshy et al. 2015). In this work, we perform an in-depth, multilingual analysis of LLM cultural alignment across several Asian nations, using religion as a critical lens. We aim to answer the following research questions:
@@
We adapt the methodology of (Santurkar et al. 2023) to answer the aforementioned questions, measuring alignment through a quantitative “representativeness” metric, based on the divergence between the model’s logit-induced probability distribution and nationally representative survey data.
Jensen-Shannon Divergence (Lin 1991) and Hellinger Distance (Hellinger 1909) are used as our primary evaluative metrics to conduct a robust and multi-faceted analysis, enabling us to pinpoint specific linguistic and demographic biases. Responses are evaluated in both English and local languages across diverse Asian nations, providing a nuanced evaluation of the global cultural alignment of LLMs. Figure 1 provides a summary of our methodology.
@@
To demonstrate how high-level distributional gaps manifest as concrete representational harms in downstream tasks, we evaluate the models using a suite of culturally aware bias benchmarks that offer broad geographic and typological coverage: CrowS-Pairs (Nangia et al. 2020), IndiBias (Sahoo et al. 2024), ThaiCLI (Kim et al. 2025), and KoBBQ (Jin et al. 2024).
Our evaluation reveals that while the contemporary models are generally representative of different Asian populations, they consistently struggle to generate true representative opinions involving religion and identity-related topics. At the same time, in our bias benchmarks, LLMs consistently rate negative framings of religious communities, such as Sunni and Shia Muslims, as more plausible than positive ones. This pattern likely reflects both uneven community representation and the influence of negative stereotypes embedded in the online discourse.
Our work underscores the need to systematically evaluate how AI models represent religious and cultural identities worldwide before their widespread adoption. To enable further research in this direction, we have made our codebase and other resources publicly available on GitHub [^1] https://github.com/HariShankar08/LLMOpinions/tree/new-method.
@@
Numerous studies demonstrate that Large Language Models (LLMs) often reflect the cultural values of English-speaking and Protestant European nations (Tao et al. 2024; Huang et al. 2023). This has led to models frequently aligning more closely with United States-centric viewpoints and failing to capture community-specific knowledge (Sukiennik et al. 2025; Etxaniz et al. 2024). Recent comparative audits further show that LLMs manifest regionally variable degrees of alignment, with notable misrepresentations persisting in Asian, African, and Latin American contexts (Bentley et al. 2025; AlKhamissi et al. 2024). Complementary work has also used LLMs to study how users engage with harmful or misleading content at scale, for example, classifying whether audiences express support or skepticism toward mental-health misinformation and revealing platform-specific amplification patterns and annotation reliability gaps (Nguyen et al. 2025).
@@
The lack of culturally representative data can lead to large gaps in societal and religious viewpoints, a particularly critical issue in multilingual and plural societies (Qin et al. 2025; Gamboa et al. 2024). A cross-lingual evaluation by (Plaza-del-Arco et al. 2024) found persistent religious stereotyping and refusals among LLMs, particularly for minority faith groups, underscoring the scarcity of systematic approaches to religion-focused NLP bias.
@@
However, measuring and mitigating these biases can be challenging. Alignment scores can flip entirely based on methodological choices like prompt formatting and question selection (Khan et al. 2025). Popular alignment techniques like Reinforcement Learning from Human Feedback often perpetuate existing biases from base models, including those related to gender (Ovalle et al. 2024; Zhang et al. 2024). LLM-based judges sometimes favor reward style over actual accuracy (Feuer et al. 2025), while using LLMs as annotators introduces systematic labeling biases that flow into downstream systems. In hate speech detection, for example, LLM-generated labels show demographic and dialect-linked disparities that prompting and ensembling strategies don’t fully address (Okpala and Cheng 2025). Studies comparing human and LLM annotators find their bias profiles differ substantially, with LLMs potentially amplifying under-detection for minority targets (Giorgi et al. 2025).
@@
Various strategies have been proposed to improve cultural alignment. Simple interventions like local-language prompting and demographic priming show both promise and clear limits in reducing bias (AlKhamissi et al. 2024; Bentley et al. 2025). Data-centric approaches use LLMs to generate semantic augmentations, such as denoising rewrites or contextual explanations, that strengthen small harmful-content datasets and improve detection even in low-resource settings (Meguellati et al. 2025). More fundamental methods involve pre-training on targeted local data to help models acquire specific cultural knowledge (Etxaniz et al. 2024). Some techniques work at deeper levels, like D2O which uses human-labeled negative examples during training (Duan et al. 2024), or FairSteer which applies corrective adjustments to model activations at inference time without retraining (Li et al. 2025). Despite this progress, the limits of these methods for comprehensive cultural adaptation remain unclear (Liu et al. 2025; Qin et al. 2025), and researchers continue developing new metrics to measure representational harms more precisely (Shin et al. 2024; Hida et al. 2024).
@@
A synthesis of current approaches suggests that scalable, region-specific audits and the curation of native survey data are necessary to ensure LLMs are deployed equitably worldwide  (Qin et al. 2025; Plaza-del-Arco et al. 2024; Bentley et al. 2025). Foundational work by  (Santurkar et al. 2023) evaluates whether model outputs reflect nationally representative opinion data, providing a methodology for such audits. Our research advances this paradigm by introducing multilinguality into existing datasets and extending evaluation beyond Western-centric benchmarks. Specifically, we augment standard resources with data in multiple languages and systematically test LLMs on tasks that foreground both religion and multilingual alignment, with particular emphasis on India and East/Southeast Asia. Leveraging large-scale, nationally representative Pew surveys and regionally salient cultural datasets (Maguire 2017), we address a critical gap: evaluating how LLMs align with local public opinion on religion across diverse linguistic contexts, especially in societies where religion remains deeply intertwined with social and political identity.
@@
While the original surveys were administered in local languages, the publicly available metadata provides the survey questions and response options only in English. To effectively test the multilingual capabilities of LLMs and align our prompts with the original survey context, a comprehensive translation of the survey instruments was required. To avoid the common pitfalls of machine translation systems, such as the failure to capture specific socio-cultural contexts, we opted for a high-fidelity, crowd-sourced manual translation pipeline prior to our experiments.
Experienced translators were recruited for each target language via crowdsourcing to ensure semantic accuracy and cultural relevance. Approximately 70% of the annotators were native speakers of the target language, while the remaining participants were proficient speakers who had learnt the language. Translators were selected based on their prior work experience in similar translation tasks, and were instructed to produce faithful, human-like translations that preserved the intent of the original survey questions and response options. All translators were paid in accordance with their preferred compensation rates, with a total cost incurred of USD 125. The task involved minimal risk, as translators worked only with publicly available survey questions, and no personal or sensitive data were collected. Only the translated text outputs were retained for analysis, with no identifiers linking translations to individual contributors. The specific languages targeted for each country are detailed in Table 1.
@@
To comprehensively evaluate the models’ representativeness and misrepresentation, we use four complementary, culturally-aware bias benchmarks: CrowS-Pairs  (Nangia et al. 2020), IndiBias  (Sahoo et al. 2024), ThaiCLI  (UpstageAI 2025; Kim et al. 2025), and KoBBQ  (Jin et al. 2024). Together, these corpora allow (i) measuring pairwise stereotyping tendencies in both masked and generative language, (ii) targeted probing of representational harms in Indian contexts (English/Hindi), (iii) evaluating Thai cultural/pragmatic alignment, and (iv) QA-style bias assessment in Korean. These benchmarks provide broader geographic and typological coverage, and enable cross-cultural comparison of representational performance. We provide a summary of the same in Table 2.
@@
IndiBias is a benchmark designed specifically for the South Asian context. Sahoo et al. (2024) is uniquely designed to test for biases along India-relevant identity axes (e.g., religion, caste, region, gender, occupation) in both English and Hindi. While smaller than some Western-centric corpora, IndiBias fills a crucial gap by explicitly testing representational harms related to South Asian identities and evaluating multilingual model behaviour in this domain.
@@
Gemini-2.5-FlashGPT-4o-Mini0505010010094.6394.6395.1795.1789.8989.8990.1790.1796.5796.5798.4198.41Representativeness score (%)Non-religiousReligiousPost-Steering
@@
0.40.40.60.60.80.811 Gemma-3-12B (SL: Sinhala vs EN) Llama-3.2-1B (Tw: ZH vs EN) Mistral-7B (Kr: KO vs EN) Jensen–Shannon Distance (lower is better)Local-language promptEnglish prompt
@@
enhijakovizh02020404060608080100100% (or count)GPT-4o biasGemini biasGPT-4o anti-stereoGemini anti-stereo
@@
20-2015-1510-105-505510101515202025253030 Shia Sunni Jain Parsi Buddhist Bahai Hindu Sufi Sikh Christian Bohra Muslim $\Delta\!>\!0$ (neg. >> pos.) $\Delta\!<\!0$ (pos. >> neg.) $\Delta\mathrm{ELO}$ (higher = more misrep.)
@@ Discussion
While this study centers on prompt-based steering for evaluating cultural alignment, deeper interventions from recent literature offer promising avenues for more profound model adaptation. Activation engineering, such as Activation Addition, enables inference-time steering by adding vectors derived from contrasting activations (e.g., positive vs. negative sentiment prompts), achieving state-of-the-art control over outputs like toxicity reduction without retraining (Turner et al. 2023). Representation engineering further refines this by mean-centring steering vectors to enhance steerability across tasks, including genre shifts or function triggering, as demonstrated in benchmarks on models like LLaMA (Zou et al. 2024; Jorgensen et al. 2023). Feedback-driven approaches, including RLHF or DPO, have been shown to amplify instruction-following while embedding local norms, though they risk overfitting or cross-cultural interference without diverse data (Sharma and others 2024a, b).
These alternatives hold potential to reshape models’ internal representations for robust handling of cultural and religious diversity, surpassing prompt-level guidance. However, their use is often constrained in production settings. Leading models like GPT-4o and Gemini-2.5-Flash operate as black-box APIs, denying access to weights or activations, and thereby making prompt engineering the primary lever for most users and applications. Thus, emphasizing prompting provides a pragmatic assessment of publicly available tools, underscoring that true deep alignment demands shifts in model training and access paradigms (Guo and others 2025).

### 23 Mar 2026 v2

@@ Abstract
Warning: This paper contains content that may be potentially offensive or upsetting.
@@ Introduction
Large language models (LLMs) have become essential tools for accessing information and generating content, with platforms such as ChatGPT handling billions of prompts from a global user base (Backlinko 2025; TechCrunch 2025). As of December 2025, ChatGPT was ranked 5th among the world’s most visited websites according to Similarweb (Similarweb 2025). Additionally, on social media platforms such as LinkedIn, recent surveys suggest that over 50% of long-form posts may be written or influenced by generative AI tools (Elad 2025). Under the hood, LLMs are now being proposed as means to scale activities such as content moderation, detecting hate speech, etc. (Kumar, Yousef, and Durumeric 2024; Singh, Bhattacharjee, and Chakraborty 2025). However, this widespread adoption comes with critical challenges. The probabilistic nature of LLMs leads to models preferentially generating viewpoints that are highly represented, and consequently, a biased world-view is derived from its training corpus (Bender et al. 2021a; Seth et al. 2025). Since Internet corpora are heavily skewed toward English, models may disproportionately reflect Western cultural sensibilities (Joshi et al. 2020). This risks marginalizing non-Western perspectives and also may lead to the dissemination of harmful stereotypes, such as linking specific religions to violence (Abid, Farooqi, and Zou 2021). As these models are increasingly integrated into education, research, and other everyday tasks, their potential to shape public discourse in ways that reinforce existing prejudices becomes a significant concern (Weidinger, Mellor, and et al. 2021; Bender et al. 2021b; Santurkar et al. 2023).
@@
While efforts to mitigate these linguistic and cultural biases are ongoing, research on cultural alignment has largely centred on American citizens and has been conducted almost exclusively in English (Santurkar et al. 2023; Durmus et al. 2023). This approach not only overlooks the majority of the world’s population but also ignores the fact that an LLM’s responses can vary significantly depending on the language of the prompt (Kang and Kim 2025). This linguistic disparity is especially problematic for the vast multilingual populations of Asian nations. For instance, while religion’s role has declined in many Western nations (Pew Research Center 2025, 2018; Australian Bureau of Statistics 2022), it remains a central and politically significant aspect of society across much of Asia (The Print 2025; Elvia Muthiariny 2024; Pew Research Center 2023). For billions of multilingual and non-English speaking users, ensuring that LLMs are culturally and linguistically representative is a critical challenge that must be addressed (Santurkar et al. 2023; Green et al. 2023).
@@
Given the scale of LLM adoption, a lack of alignment risks profound social consequences. These risks are already evident on social media, where the proliferation of AI-generated content has contributed to polarised discourse and marginalised certain groups, such as the LGBTQ+ community (Kerwin 2024; Bakshy, Messing, and Adamic 2015). In this work, we perform an in-depth, multilingual analysis of LLM cultural alignment across several Asian nations, using religion as a critical lens. We aim to answer the following research questions:
@@
We adapt the methodology of Santurkar et al. (2023) to answer the aforementioned questions, measuring alignment through a quantitative “representativeness” metric, based on the divergence between the model’s logit-induced probability distribution and nationally representative survey data. We use Jensen-Shannon Divergence (Lin 1991) and Hellinger Distance (Hellinger 1909) as our primary evaluative metrics to conduct a robust and multi-faceted analysis, enabling us to pinpoint specific linguistic and demographic biases. Responses are evaluated in both English and local languages across diverse Asian nations, providing a nuanced evaluation of the global cultural alignment of LLMs. Figure 1 provides a summary of our methodology.
@@
To demonstrate how high-level distributional gaps manifest as concrete representational harms in downstream tasks, we evaluate the models using a suite of culturally aware bias benchmarks that offer broad geographic and typological coverage: CrowS-Pairs (Nangia et al. 2020), IndiBias (Sahoo et al. 2024), ThaiCLI (Kim et al. 2025), and KoBBQ (Jin et al. 2024). Our evaluation reveals that while the contemporary models are generally representative of different Asian populations, they consistently struggle to generate true representative opinions involving religion and identity-related topics. At the same time, in our bias benchmarks, LLMs consistently rate negative framings of religious communities, such as Sunni and Shia Muslims, as more plausible than positive ones. This pattern likely reflects both uneven community representation and the influence of negative stereotypes embedded in the online discourse.
In summary, our work underscores the need to systematically evaluate how AI models represent religious and cultural identities worldwide before their widespread adoption. To enable further research in this direction, we have made our codebase and other resources publicly available on GitHub [^1] https://github.com/HariShankar08/LLMOpinions.
@@
Numerous studies demonstrate that Large Language Models (LLMs) often reflect the cultural values of English-speaking and Protestant European nations (Tao et al. 2024; Huang et al. 2023). This has led to models frequently aligning more closely with United States-centric viewpoints and failing to capture community-specific knowledge (Sukiennik et al. 2025; Etxaniz et al. 2024). Recent comparative audits further show that LLMs manifest regionally variable degrees of alignment, with notable misrepresentations persisting in Asian, African, and Latin American contexts (Bentley, Evans, and Bull 2025; AlKhamissi et al. 2024). Complementary work has also used LLMs to study how users engage with harmful or misleading content at scale, for example, classifying whether audiences express support or skepticism toward mental-health misinformation and revealing platform-specific amplification patterns and annotation reliability gaps (Nguyen et al. 2025).
@@
The lack of culturally representative data can lead to large gaps in societal and religious viewpoints, a particularly critical issue in multilingual and plural societies (Qin et al. 2025; Gamboa, Feng, and Lee 2024; Chhikara, Kumar, and Chakraborty 2025). A cross-lingual evaluation by (del Arco, Pelloni, and Zampieri 2024) found persistent religious stereotyping and refusals among LLMs, particularly for minority faith groups, underscoring the scarcity of systematic approaches to religion-focused NLP bias.
@@
However, measuring and mitigating these biases can be challenging. Alignment scores can flip entirely based on methodological choices like prompt formatting and question selection (Khan, Casper, and Hadfield-Menell 2025). Popular alignment techniques like Reinforcement Learning from Human Feedback often perpetuate existing biases from base models, including those related to gender (Ovalle et al. 2024; Zhang et al. 2024). LLM-based judges sometimes favor reward style over actual accuracy (Feuer et al. 2025), while using LLMs as annotators introduces systematic labeling biases that flow into downstream systems. In hate speech detection, for example, LLM-generated labels show demographic and dialect-linked disparities that prompting and ensembling strategies don’t fully address (Okpala and Cheng 2025). Studies comparing human and LLM annotators find their bias profiles differ substantially, with LLMs potentially amplifying under-detection for minority targets (Giorgi et al. 2025).
@@
Various strategies have been proposed to improve cultural alignment. Simple interventions like local-language prompting and demographic priming show both promise and clear limits in reducing bias (AlKhamissi et al. 2024; Bentley, Evans, and Bull 2025; Chhikara et al. 2024). Data-centric approaches use LLMs to generate semantic augmentations, such as denoising rewrites or contextual explanations, that strengthen small harmful-content datasets and improve detection even in low-resource settings (Meguellati et al. 2025). More fundamental methods involve pre-training on targeted local data to help models acquire specific cultural knowledge (Etxaniz et al. 2024). Some techniques work at deeper levels, like D2O which uses human-labeled negative examples during training (Duan et al. 2024), or FairSteer which applies corrective adjustments to model activations at inference time without retraining (Li et al. 2025). Despite this progress, the limits of these methods for comprehensive cultural adaptation remain unclear (Liu, Korhonen, and Gurevych 2025; Qin et al. 2025), and researchers continue developing new metrics to measure representational harms more precisely (Shin et al. 2024; Hida, Yamaguchi, and Hanawa 2024).
@@
A synthesis of current approaches suggests that scalable, region-specific audits and the curation of native survey data are necessary to ensure LLMs are deployed equitably worldwide  (Qin et al. 2025; del Arco, Pelloni, and Zampieri 2024; Bentley, Evans, and Bull 2025). Foundational work by  (Santurkar et al. 2023) evaluates whether model outputs reflect nationally representative opinion data, providing a methodology for such audits. Our research advances this paradigm by introducing multilinguality into existing datasets and extending evaluation beyond Western-centric benchmarks. Specifically, we augment standard resources with data in multiple languages and systematically test LLMs on tasks that foreground both religion and multilingual alignment, with particular emphasis on India and East/Southeast Asia. Leveraging large-scale, nationally representative Pew surveys and regionally salient cultural datasets (Maguire 2017), we address a critical gap: evaluating how LLMs align with local public opinion on religion across diverse linguistic contexts, especially in societies where religion remains deeply intertwined with social and political identity.
@@
While the original surveys were administered in local languages, the publicly available metadata provides the survey questions and response options only in English. To effectively test the multilingual capabilities of LLMs and align our prompts with the original survey context, a comprehensive translation of the survey instruments was required. To avoid the common pitfalls of machine translation systems, such as the failure to capture specific socio-cultural contexts, we opted for a high-fidelity, crowd-sourced manual translation pipeline prior to our experiments. Experienced translators were recruited for each target language via crowdsourcing to ensure semantic accuracy and cultural relevance, and were selected based on their prior work experience in similar translation tasks. Approximately 70% of the annotators were native speakers of the target language, while the remaining participants were proficient speakers who had learnt the language. To ensure high-quality outputs, translators were explicitly instructed to:
  1. 1.
    Preserve the original meaning and intent of the survey questions and response options.
  2. 2.
    Maintain cultural nuance and context, especially concerning sensitive topics.
  3. 3.
    Ensure the resulting translations sound natural and as human-like as possible.
To evaluate the reliability of the translation process and ensure consistency, overlapping subsets of the survey questions were assigned to multiple annotators to measure inter-annotator agreement, where we noted a strong average agreement score (Cohen’s $\kappa=0.82$). Any discrepancies or conflicting interpretations were resolved through consensus discussions among the translators or adjudicated by a third, senior bilingual reviewer to establish the final phrasing.
As an additional validation measure, a subset of the translated survey questions was back-translated into English using machine translation tools and compared against the original text to identify potential inconsistencies. We ensured that important terms (e.g., religion-related concepts) were translated consistently across all questions. Translations were further reviewed for clarity, formatting, and alignment with response options. Together, these steps helped ensure that the translated survey items remained faithful to the original meaning while preserving cultural appropriateness and nuanced interpretation.
All translators were paid in accordance with their preferred compensation rates, with a total cost incurred of USD 125. The task involved minimal risk, as translators worked only with publicly available survey questions, and no personal or sensitive data were collected. Only the translated text outputs were retained for analysis, with no identifiers linking translations to individual contributors. The specific languages targeted for each country are detailed in Table 1.
@@
To comprehensively evaluate the models’ representativeness and misrepresentation, we use four complementary, culturally-aware bias benchmarks: CrowS-Pairs  (Nangia et al. 2020), IndiBias  (Sahoo et al. 2024), ThaiCLI  (UpstageAI 2025; Kim et al. 2025), and KoBBQ  (Jin et al. 2024). Together, these corpora allow (i) measuring pairwise stereotyping tendencies in both masked and generative language, (ii) targeted probing of representational harms in Indian contexts (English/Hindi), (iii) evaluating Thai cultural/pragmatic alignment, and (iv) QA-style bias assessment in Korean. These benchmarks provide broader geographic and typological coverage and enable cross-cultural comparison of representational performance. We provide a summary of the same in Table 2.
@@
IndiBias is a benchmark designed specifically for the South Asian context (Sahoo et al. 2024), which is uniquely designed to test for biases along India-relevant identity axes (e.g., religion, caste, region, gender, occupation) in both English and Hindi. While smaller than some Western-centric corpora, IndiBias fills a crucial gap by explicitly testing representational harms related to South Asian identities and evaluating multilingual model behaviour in this domain.
@@ Discussion
While this study centers on prompt-based steering for evaluating cultural alignment, deeper interventions from recent literature offer promising avenues for more profound model adaptation. Activation engineering, such as Activation Addition, enables inference-time steering by adding vectors derived from contrasting activations (e.g., positive vs. negative sentiment prompts), achieving state-of-the-art control over outputs like toxicity reduction without retraining (Turner et al. 2023). Representation engineering further refines this by mean-centring steering vectors to enhance steerability across tasks, including genre shifts or function triggering, as demonstrated in benchmarks on models like LLaMA (Zou et al. 2024; Jorgensen et al. 2023). Feedback-driven approaches, including RLHF or DPO, have been shown to amplify instruction-following while embedding local norms, though they risk overfitting or cross-cultural interference without diverse data (Sharma et al. 2024a, b).
These alternatives hold potential to reshape models’ internal representations for robust handling of cultural and religious diversity, surpassing prompt-level guidance. However, their use is often constrained in production settings. Leading models like GPT-4o and Gemini-2.5-Flash operate as black-box APIs, denying access to weights or activations, and thereby making prompt engineering the primary lever for most users and applications. Thus, emphasizing prompting provides a pragmatic assessment of publicly available tools, underscoring that true deep alignment demands shifts in model training and access paradigms (Guo et al. 2025).
@@
Acknowledgements
The authors gratefully acknowledge the financial support provided by the EkStep Foundation. We also thank the members of the Precog Research Group of IIIT Hyderabad for their help and guidance during the experimental design phase. Finally, we thank the anonymous reviewers whose feedback substantially improved the quality of the paper.